The function f(x, y) ={x²y/x⁴ +y² , (x, y) ≠0 and 0 , (x, y) =0 }is not continuous at (0,0).


Say true or false.


1
Expert's answer
2022-01-20T10:52:37-0500

Evaluate "\lim\limits_{(x,y)\to(0,0)}f(x, y)" along path "y=mx"


"\lim\limits_{(x,mx)\to(0,0)}\dfrac{x^2(mx)}{x^4+(mx)^2}=\lim\limits_{(x,mx)\to(0,0)}\dfrac{mx}{x^2+m^2}=0"

Evaluate "\lim\limits_{(x,y)\to(0,0)}f(x, y)" along path "y=kx^2"


"\lim\limits_{(x,kx^2)\to(0,0)}\dfrac{x^2(kx^2)}{x^4+(kx^2)^2}=\lim\limits_{(x,kx^2)\to(0,0)}\dfrac{k}{1+k^2}=\dfrac{k}{1+k^2}\not=0"

Therefore "\lim\limits_{(x,y)\to(0,0)}f(x, y)" does not exist.

Therefore the function "f(x, y)" is not continuous at "(0,0)."

The statement is False.